OpenStudy (anonymous):

Divide: (7 + 4i) ÷ (2 + 3i)

OpenStudy (anonymous):

wait i got it i was doing something wrong

OpenStudy (anonymous):

multiply by the conjugate of the denominator, and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (anonymous):

the denominator you do in your head, \(2^2+3^2=4+9=13\) the numerator you have to find \((7+4i)(2-3i)\)

OpenStudy (anonymous):

(2-3i) i have to multi[ly to bottom and top

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah thats what i meesed up on